List of publications of Endre Csáki

[1] Csáki, E.: On two modifications of the Wilcoxon test. Publ. Math. Inst. Hung. Acad. Sci. 4 (1959), 313–319.

[2] Csáki, E. and Vincze, I.: On some problems connected with the Galton test. Publ. Math. Inst. Hung. Acad. Sci. 6 (1961), 97–109.

[3] Csáki, E.: On the number of intersections in the one–dimensional random walk. Publ. Math. Inst. Hung. Acad. Sci. 6 (1961), 281–286.

[4] Csáki, E. and Németh, E.: Methods of mathematical statistics for evaluating electric breakdown measuring series. Periodica Polyt. Electr. Eng. 7 (1963), 9–35.

[5] Csáki, E. and Vincze, I.: On some combinatorial relations concerning the symmetric random walk. Acta Sci. Math. (Szeged) 24 (1963), 231–235.

[6] Csáki, E. and Vincze, I.: Two joint distribution laws in the theory of order statistics. Mathematica (Cluj) 5 (1963), 27–37.

[7] Csáki, E. and Vincze, I.: On some distributions connected with the arcsine law. Publ. Math. Inst. Hung. Acad. Sci. 8 (1963), 281–291.

[8] Csáki, E., Balogh, A. and Sarkadi, K.: Statistical methods of reliability. (In Hungarian.) Hiradástechnika 15 (1964), 332–339.

[9] Csáki, E.: Nonparametric methods. Publ. Inst. Autom. Hung. Acad. Sci. 11 (1965), 188–202.

[10] Csáki, E.: On the statistical problems of life tests of electrical components. Wissl. Zeitschrift der Humboldt Univ. 16 (1967), 6–8.

[11] Csáki, E.: An iterated logarithm law for semimartingales and its application to empirical distribution function. Studia Sci. Math. Hungar. 3 (1968), 287–292.

[12] Csáki, E.: On the number of intersections between two empirical distribution functions. In: Studies in Math. Statist. Publ House Acad. Sci. Hungar., Budapest, 1968, pp. 27–32.

[13] Csáki, E. and Tusnády, G.: On the number of intersections and the ballot theorem. Periodica Math. Hungar. 2 (1972), 5–13.

[14] Csáki, E.: Some notes on the law of the iterated logarithm for empirical distribution function Colloquia Math. Soc. J. Bolyai 11 (1974), 47–58.

[15] Csáki, E.: Investigations concerning the empirical distribution function. (In Hungarian.) Magyar Tud. Akad. Mat. Fiz. Oszt. Közl. 22 (1977), 239–327. Translation: Selected Transl. Math. Statist. Probab. 15 (1981), 229–317.

[16] Csáki, E. and Vincze, I.: A lemma on real functions and its application in the nonparametric theory. In: Proc. Symp. to honour J. Neyman, Warsaw, 1977, pp. 71–80.

[17] Csáki, E.: The law of the iterated logarithm for normalized empirical distribution function. Z. Wahrscheinlichkeitstheorie verw. Gebiete 38 (1977), 147–167.

[18] Csáki, E. and Vincze, I.: On modified forms of Pearson’s chi square. In: Proc. ISI Conference, Delhi, 1977, pp. 669–672.

[19] Csáki, E.: On some distributions concerning maximum and minimum of a Wiener process. Colloquia Math. Soc. J. Bolyai 21 (1977), 43–52.

[20] Csáki, E.: On the lower limits of maxima and minima of Wiener process and partial sums. Z. Wahrscheinlichkeitstheorie verw. Gebiete 43 (1978), 205–221.

[21] Csáki, E. and Vincze, I.: On the limiting distribution laws of statistics analogous to Pearson’s chi–square. Math. Operationsforsch. Statist. Ser. Statistics 9 (1978), 531–548.

[22] Csáki, E. and Révész, P.: How big must be the increments of a Wiener process? Acta Math. Acad. Sci. Hungar. 33 (1979), 37–49.

[23] Csáki, E. and Földes, A.: On the first empty cell. Studia Sci. Math. Hungar. 11 (1976), 373–382.

[24] Csáki, E. and Mohanty, S.G.: Excursion and meander in random walk. Canad. J. Statist. 9 (1981), 57–70.

[25] Csáki, E.: A relation between Chung’s and Strassen’s laws of the iterated logarithm. Z. Wahrscheinlichkeitstheorie verw. Gebiete 54 (1980), 287–301.

[26] Csáki, E.: Iterated logarithm laws for the square integral of a Wiener process. Lecture Notes in Statist. Springer, New York, 1981, pp. 42–53.

[27] Csáki, E.: On the standardized empirical distribution function. Colloquia Math. Soc. J. Bolyai 32 (1980), 123–138.

[28] Csáki, E.: A note on the lower limit of the standardized empirical distribution function. Mimo. Ser. No. 81-19, 1981, Dept. Statistics, Purdue Univ.

[29] Csáki, E. and Révész, P.: Strong invariance for local times. Z. Wahrscheinlichkeitstheorie verw. Gebiete 62 (1983), 263–278.

[30] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: How big are the increments of the local time of a Wiener process? Ann. Probab. 11 (1983), 593–608.

[31] Csáki, E. and Révész, P.: A combinatorial proof of a theorem of P. Lévy on the local time. Acta Sci. Math. (Szeged) 45 (1983), 119–129.

[32] Csáki, E.: Empirical distribution function. In: Nonparametric Methods, Handbook of Statistics, Vol. 4, 1984, pp. 405–430.

[33] Csáki, E. and Földes, A.: On the narrowest tube of a Wiener process. Colloquia Math. Soc. J. Bolyai 36 (1982), 173–197.

[34] Csáki, E. and Földes, A.: How big are the increments of the local time of a simple symmetric random walk? Colloquia Math. Soc. J. Bolyai 36 (1984), 199–221.

[35] Csáki, E.: On small values of the square integral of a multiparameter Wiener process. In: Proc. of the 3rd Pannonian Symp. on Math. Statist., Akadémiai Kiadó, Budapest, 1983, pp. 19–26.

[36] Csáki, E. and Földes, A.: How big are the increments of the local time of a recurrent random walk? Z. Wahrscheinlichkeitstheorie verw. Gebiete 65 (1983), 307–322.

[37] Csáki, E. and Földes, A.: The narrowest tube of a recurrent random walk. Z. Wahrscheinlichkeitstheorie verw. Gebiete 66 (1984), 387–405.

[38] Csáki, E. and Révész, P.: On a theorem of Chung and Erdős. (In Hungarian.) Mat. Lapok 31 (1978–1983), 243–249.

[39] Csáki, E., Erdős, P. and Révész, P.: On the length of the longest excursion. Z. Wahrscheinlichkeitstheorie verw. Gebiete 68 (1985), 365–382.

[40] Csáki, E., Földes, A. and Komlós, J.: Limit theorems for Erdős–Rényi type problems. Studia Sci. Math. Hungar. 22 (1987), 321–332.

[41] Csáki, E. and Földes, A.: How small are the increments of the local time of a Wiener process? Ann. Probab. 14 (1986), 533–546.

[42] Csáki, E. and Földes, A.: How small are the increments of the local time of a recurrent random walk? Periodica Math. Hungar. 17 (1986), 127–136.

[43] Csáki, E. and Mohanty, S.G.: Some joint distributions for conditional random walks. Canad. J. Statist. 14 (1986), 19–28.

[44] Csáki, E.: Some applications of the classical formula on ruin probabilities. J. Statist. Plann. Inference 14 (1986), 35–42.

[45] Csáki, E., Földes, A. and Révész, P.: On the maximum of a Wiener process and its location. Probab. Th. Rel. Fields 76 (1987), 477–497.

[46] Csáki, E. and Földes, A.: On the length of the longest flat interval. In: Proc. of the 5th Pannonian Symp. on Math. Statist., Visegrád, 1988, pp. 23–33.

[47] Csáki, E. and Földes, A.: A note on the stability of the local time of a Wiener process. Stoch. Process. Appl. 25 (1987), 203–213.

[48] Csáki, E., Földes, A. and Salminen, P.: On the joint distribution of the maximum and its location for a linear diffusion. Ann. Inst. H. Poincaré Probab. Statist. 23 (1987), 179–194.

[49] Csáki, E. and Földes, A.: On the local time process standardized by the local time at zero. Acta Math. Acad. Sci. Hungar. 52 (1988), 175–186.

[50] Csáki, E. and Grill, K.: On the large values of the Wiener process. Stoch. Process. Appl. 27 (1988), 43–56.

[51] Csáki, E., Földes, A. and Kasahara, Y.: Around Yor’s theorem on the Brownian sheet and local time. J. Math. Kyoto Univ. 28 (1988), 373–381.

[52] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Brownian local time approximated by a Wiener sheet. Ann. Probab. 17 (1989), 516–537.

[53] Csáki, E.: An integral test for the supremum of Wiener local time. Probab. Th. Rel. Fields 83 (1989), 207–217.

[54] Csáki, E., Mohanty, S.G. and Saran, J.: On random walks in a plane. Ars Combinatoria 29 (1990), 309–318.

[55] Csáki, E.: A lim inf result in Strassen’s law of the iterated logarithm. Colloquia Math. Soc. J. Bolyai 57 (1990), 83–93.

[56] Csáki, E. and Csörgő, M.: Fernique type inequalities for not necessarily Gaussian processes. C.R. Math. Rep. Acad. Sci. Canada 12 (1990), 149–154.

[57] Csáki, E.: Lim inf results for the standardized empirical distribution function. In: Prob. Theory and Math. Stat., Vol. I, B. Grigelionis et al. (Eds.), 1990, 238–244.

[58] Csáki, E., Csörgő, M., Lin, Z.Y. and Révész, P.: On infinite series of independent Ornstein–Uhlenbeck processes. Stoch. Proces. Appl. 39 (1991), 25–44.

[59] Csáki, E.: On the increments of additive functionals. Studia Sci. Math. Hungar. 26 (1991), 185–199.

[60] Csáki, E.: A note on ”Three problems on the random walk in Z^{d}” by P. Erdős and P. Révész. Studia Sci. Math. Hungar. 26 (1991), 201–205.

[61] Csáki, E. and Csörgő, M.: Inequalities for increments of stochastic processes and moduli of continuity. Ann. Probab. 20 (1992), 1031–1052.

[62] Csáki, E., Csörgő, M. and Shao, Q.M.: Fernique type inequalities and moduli of continuity for l^{2}-valued Ornstein–Uhlenbeck processes. Ann. Inst. Henri Poincaré 28 (1992), 479–517.

[63] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Strong approximations of additive functionals. J. Theor. Probab. 5 (1992), 679–706.

[64] Csáki, E.: A discrete Feynman–Kac formula. J. Statist. Plann. Inf. 34 (1993), 63–73.

[65] Csáki, E.: On the local time of Wiener process and random walk. Lecture Notes of 7th International Summer School on Probability Theory and Mathematical Statistics, 1993, pp. 32–59.

[66] Csáki, E., Földes, A. and Révész, P.: On almost sure local and global central limit theorems. Probab. Th. Rel. Fields 97 (1993), 321–337.

[67] Csáki, E.: Some joint distributions in Bernoulli excursions. J. Appl. Probab. 31A (1994), 239–250.

[68] Csáki, E.: Strong limit theorems for empirical processes. Recent Advances in Statistics and Probability, Proc. 4th IMSIBAC, ed. J. P. Vilaplana and M. L. Puri, VsP, Zeist, 1994, pp. 247–254.

[69] Csáki, E. and Földes, A.: On the logarithmic average of additive functionals. Statist. Probab. Lett. 22 (1995), 261–268.

[70] Csáki, E., Csörgő, M. and Shao, Q.M.: Moduli of continuity for ℓ^{p}-valued Gaussian processes. Acta Sci. Math. (Szeged) 60 (1995), 149–175.

[71] Csáki, E. and Csörgő, M.: On additive functionals of Markov chains. J. Theor. Probab. 8 (1995), 905–919.

[72] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Global Strassen-type theorems for iterated Brownian motion. Stoch. Process. Appl. 59 (1995), 321–341.

[73] Csáki, E. and Földes, A.: On the length of the longest monotone block. Studia Sci. Math. Hungar. 31 (1996), 35–46.

[74] Csáki, E. and Salminen, P.: On additive functionals of diffusion processes. Studia Sci. Math. Hungar. 31 (1996), 47–62.

[75] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: The local time of iterated Brownian motion. J. Theor. Probab. 9 (1996), 717–743.

[76] Berkes,I. and Csáki, E.: On the pointwise central limit theorem and mixtures of stable distributions. Statist. Probab. Lett. 29 (1996), 361–368.

[77] Csáki, E.: On some results for Bernoulli excursions. J. Statist. Plann. Inf. 54 (1996), 45–54.

[78] Csáki, E., Földes, A., Révész, P.: Random walk with alternating excursions. Studia Sci. Math. Hungar. 32 (1996), 267–280.

[79] Csáki, E., Földes, A. and Révész, P.: Strassen theorems for a class of iterated processes. Trans. Amer. Math. Soc. 349 (1997), 1153–1167.

[80] Csáki, E. and Földes, A.: On the logarithmic avarage of iterated processes. Statist. Probab. Lett. 33 (1997), 347–358.

[81] Csáki, E.: Some results for two-dimensional random walk. In: Advances in Combinatorial Methods and Applications to Probability and Statistics, ed. N. Balakrishnan, Birkhäuser, Boston, 1997, pp. 115–124.

[82] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: On the occupation time of an iterated process having no local time. Stoch. Process. Appl. 70 (1997), 199–217.

[83] Berkes, I., Csáki, E. and Horváth, L.: Almost sure central limit theorems under minimal conditions. Statist. Probab. Lett. 37 (1998), 67–76.

[84] Csáki, E., Révész, P. and Rosen, J.: Functional laws of the iterated logarithm for local times of recurrent random walks on Z^{2}. Ann. Inst. Henri Poincaré, Probabilités et Statistiques 34 (1998), 545–563.

[85] Csáki, E., Földes, A. and Révész, P.: A strong invariance principle for the local time difference of a simple symmetric planar random walk. Studia Sci. Math. Hungar. 34 (1998), 25–39.

[86] Csáki, E. and Shi, Z.: Large favourite sites of simple random walk and the Wiener process. Electronic J. Probab. 3 (1998), Paper No. 14, 1–31.

[87] Csáki, E. and Shi, Z.: Some liminf results for two-parameter processes. Stoch. Process. Appl. 78 (1998), 27–46.

[88] Csáki, E. and Földes, A.: On two ergodic properties of self-similar processes. In: Asymptotic Methods in Probability and Statistics, ed. B. Szyszkowicz, Elsevier, Amsterdam, 1998, pp. 97–111.

[89] Csáki, E. and Földes, A.: On asymptotic independence of partial sums. In: Asymptotic Methods in Probability and Statistics, ed. B. Szyszkowicz, Elsevier, Amsterdam, 1998, pp. 373–385.

[90] Csáki, E., König, W. and Shi, Z.: An embedding for the Kesten–Spitzer random walk in random scenery. Stoch. Process. Appl. 82 (1999), 283–292.

[91] Berkes, I. Csáki, E. and Csörgő, S.: Almost sure limit theorems for the St. Petersburg game. Statist. Probab. Lett. 45 (1999), 23–30.

[92] Csáki, E., Khoshnevisan, D. and Shi, Z.: Capacity estimates, boundary crossings and the Ornstein–Uhlenbeck process in Wiener space. Electronic Comm. Probab. 4 (1999), 103–109.

[93] Csáki, E., Földes, A., Révész, P. and Shi, Z.: On the excursions of two-dimensional random walk and Wiener process. Bolyai Society Mathematical Studies 9 (1999), 43–58.

[94] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Asymptotic properties of integral functionals of geometric stochastic processes. J. Applied Probab. 37 (2000), 480– 493.

[95] Csáki, E., Csörgő, M., Földes, A. and Shi, Z.: Increment sizes of the principal value of Brownian local time. Probab. Th. Rel. Fields 117 (2000), 515–531.

[96] Csáki, E., Khoshnevisan, D. and Shi, Z.: Boundary crossings and the distribution function of the maximum of Brownian sheet. Stoch. Process. Appl. 90 (2000), 1–18.

[97] Csáki, E., Révész, P. and Shi, Z.: Favourite sites, favourite values and jump sizes for random walk and Brownian motion. Bernoulli 6 (2000), 951–975.

[98] Csáki, E. and Földes, A.: Asymptotic independence and additive functionals. J. Theoret. Probab. 13 (2000), 1123–1144.

[99] Csáki, E. and Hu, Y.: Asymptotic properties of ranked heights in Brownian excursions. J. Theoret. Probab. 14 (2001), 77–96.

[100] Csáki, E., Révész, P. and Shi, Z.: A strong invariance principle for two-dimensional random walk in random scenery. Stoch. Process. Appl. 92 (2001), 181–200.

[101] Csáki, E., Csörgő, M., Földes, A. and Shi, Z.: Path properties of Cauchy’s principal values related to local time. Studia Sci. Math. Hungar. 38 (2001), 149–169.

[102] Berkes, I. and Csáki, E.: A universal result in almost sure central limit theory. Stoch. Process. Appl. 94 (2001), 105–134.

[103] Csáki, E., Révész, P. and Shi,Z.: Long excursions of a random walk. J. Theoret. Probab. 14 (2001), 821–844.

[104] Csáki, E.: Some results for two-dimensional random walk, II. Additive functionals. J. Statist. Plann. Inf. 101 (2002), 49–56.

[105] Csáki, E., Csörgő, M., Földes, A., Shi, Z. and Zitikis, R.: Pointwise and uniform asymptotics of the Vervaat error process. J. Theoret. Probab. 15 (2002), 845–875.

[106] Csáki, E. and Gonchigdanzan, K.: Almost sure limit theorems for the maximum of stationary Gaussian sequences. Statist. Probab. Lett. 58 (2002), 195–203.

[107] Csáki E.: Határértéktételek a lokális idő Hilbert-transzformáltjára (Limit theorems for the Hilbert transform of local time). Magyar Tudományos Akadémia, Közgyűlési Előadások, 1999-2002, 97–102.

[108] Berkes, I., Csáki, E., Csörgő, S. and Megyesi, Z.: Almost sure limit theorems for sums and maxima from the domain of geometric partial attraction of semistable laws. In: Limit Theorems in Probability and Statistics, Vol. I. (I. Berkes, E. Csáki, M. Csörgő, eds.) J. Bolyai Mathematical Society, Budapest, 2002 pp. 133–157.

[109] Csáki, E., Csörgő, M., Földes, A. and Shi, Z.: On a class of additive functionals of two-dimensional Brownian motion and random walk. In: Limit Theorems in Probability and Statistics, Vol. I. (I. Berkes, E. Csáki, M. Csörgő, eds.) J. Bolyai Mathematical Society, Budapest, 2002, pp. 321–345.

[110] Csáki, E. and Hu, Y.: On the joint asymptotic behaviours of ranked heights of Brownian excursions. In: Limit Theorems in Probability and Statistics, Vol. I. (I. Berkes, E. Csáki, M. Csörgő, eds.) J. Bolyai Mathematical Society, Budapest, 2002, pp. 347–363.

[111] Csáki, E., Shi, Z. and Yor, M.: Fractional Brownian motions as “higher-order” fractional derivatives of Brownian local times. In: Limit Theorems in Probability and Statistics, Vol. I. (I. Berkes, E. Csáki, M. Csörgő, eds.) J. Bolyai Mathematical Society, Budapest, 2002, pp. 365–387.

[112] Csáki, E., Földes, A. and Shi, Z.: A joint functional law for the Wiener process and principal value. Studia Sci. Math. Hungar. 40 (2003), 213–241.

[113] Csáki, E. and Hu, Y.: Lengths and heights of random walk excursions. In: Discrete Mathematics and Theoretical Computer Science, Discrete Random Walks, DRW03, (C. Banderier, C. Krattenthaler, eds.) Conference Volume AC, 2003, pp. 45–52.

[114] Csáki, E. and Hu, Y.: Invariance principles for ranked excursion lengths and heights. Electronic Comm. Probab. 9 (2004), 14–21.

[115] Csáki, E., Földes, A. and Hu, Y.: Strong approximations of additive functionals of a planar Brownian motion. Stoch. Process. Appl. 109 (2004), 263–293.

[116] Csáki, E. and Hu, Y.: On the ranked excursion heights of a Kiefer process. J. Theoret. Probab. 17 (2004), 145–163.

[117] Csáki, E. Révész, P. and Shi, Z.: Large void zones and occupation times for coalescing random walks. Stoch. Process. Appl. 111 (2004), 97–118.

[118] Csáki, E., Földes, A. and Shi, Z.: Our joint work with Miklós Csörgő. In: Asymptotic Methods in Stochastics. Festschrift for Miklós Csörgő. (L. Horváth, B. Szyszkowicz, eds.) Fields Institute Communications, Vol. 44, American Mathematical Society, 2004, pp. 3–24.

[119] Csáki, E.: István Vincze (1912–1999) and his contribution to lattice path combinatorics and statistics. J. Statist. Plann. Inf. 135 (2005), 3–17.

[120] Csáki, E. and Hu, Y.: On the increments of the principal value of Brownian local time. Electronic J. Probab. 10 (2005), 925–947.

[121] Csáki, E., Földes, A., Révész, P., Rosen, J. and Shi, Z.: Frequently visited sets for random walks. Stoch. Process. Appl. 115 (2005), 1503–1517.

[122] Csáki, E., Földes, A. and Révész, P.: Maximal local time of a d-dimensional random walk on subsets. J. Theoret. Probab. 18 (2005), 687–717.

[123] Csáki, E.: Mathematical Statistics. In: A Panorama of Hungarian Mathematics in the Twentieth Century. (J. Horváth, ed.) Bolyai Society Mathematical Studies, Vol. 14, Springer, Berlin and J. Bolyai Society, Budapest, 2005, pp. 491–521.

[124] Csáki, E., Földes, A. and Révész, P.: Heavy points of a d-dimensional simple random walk. Statist. Probab. Lett. 76 (2006), 45–57.

[125] Csáki, E., Csörgő, M., Rychlik, Z. and Steinebach, J.: On Vervaat and Vervaaterror-type processes for partial sums and renewals. J. Statist. Plann. Inf. 137 (2007), 953–966.

[126] Csáki, E. and Hu, Y.: Strong approximations of three-dimensional Wiener sausages. Acta Math. Hungar. 114 (2007), 205–226.

[127] Csáki, E., Földes, A. and Révész, P.: On the local times of transient random walks. Acta Appl. Math. 96 (2007), 147–158.

[128] Csáki, E., Földes, A. and Révész, P.: Joint asymptotiv behavior of local and occupation times of random walk in higher dimension. Studia Sci. Math. Hungar. 44 (2007), 535–563.

[129] Csáki, E., Földes, A. and Révész, P.: On the behavior of random walk around heavy points. J. Theoret. Probab. 20 (2007), 1041–1057.

[130] Csáki, E., Földes, A. and Révész, P.: On the local time of the asymmetric Bernoulli walk. Acta Sci. Math. (Szeged) 74 (2008), 349–379.

[131] Csáki, E., Földes, A. and Révész, P.: Transient nearest neighbor random walk on the line. J. Theoret. Probab. 22 (2009), 100–122.

[132] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Strong limit theorems for a simple random walk on the 2-dimensional comb. Electr. J. Probab. 14 (2009), 2371–2390.

[133] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Random walk local time approximated by a Brownian sheet combined with an independent Brownian motion. Ann. Inst. H. Poincaré, Probab. Statist. 45 (2009), 515–544.

[134] Csáki, E., Földes, A. and Révész, P.: Transient nearest neighbor random walk and Bessel process. J. Theoret. Probab. 22 (2009), 992–1009.

[135] Csáki, E., Földes, A. and Révész, P.: On the number of cutpoints of the transient nearest neighbor random walk on the line. J. Theoret. Probab. 23 (2010), 624– 638.

[136] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: On the supremum of iterated local time. Publ. Math. Debrecen 76 (2010), 255–270.

[137] Csáki, E., Csörgő, M. and Kulik, R.: On Vervaat processes for sums and renewals in weakly dependent cases. Dependence in Probability, Analysis and Number Theory. A Volume in Memory of Walter Philipp, edited by I. Berkes, R.C. Bradley, H. Dehling, M. Peligrad, R. Tichy. Kendrick Press, 2010, pp. 145–156.

[138] Csáki, E., Csörgő, M. Földes, A. and Révész, P.: On the local time of random walk on the 2-dimensional comb. Stoch. Process. Appl. 121 (2011), 1290–1314.

[139] Csáki, E., Csörgő, M. Földes, A. and Révész, P.: Random walk on half-plane half-comb structure. Ann. Mathematicae et Informaticae 39 (2012), 29–44.

[140] Csáki, E., Csörgő, M. Földes, A. and Révész, P.: Strong limit theorems for anisotropic random walks on Z^{2}. Periodica Math. Hungar. 67 (2013) 71–94.

[141] Csáki, E., Földes, A. and Révész, P.: Some results and problems for anisotropic random walk on the plane. In: Asymptotic Laws and Methods in Stochastics: A Volume in Honour of Miklós Csörgő. Fields Institute Communications 76, Springer, New York (2015) 55–75.

[142] Csáki, E. and Csörgő, M.: On Bahadur-Kiefer type processes for sums and renewals in dependent cases. In: Mathematical Statistics and Limit Theorems: Conference in Honor of Paul Deheuvels. Springer, New York (2015) 93–103.

[143] Csáki, E., Csörgő, M. and Kulik, R.: Strong approximations for long memory sequences based partial sums, counting and their Vervaat processes. Periodica Math. Hungar. 73 (2016) 208–223.

[144] Csáki, E., Csörgő, M., Földes, A. and Révész, P.: Some limit theorems for heights of random walks on a spider. J. Theoret. Probab. 29 (2016) 1685–1709.

[145] Csáki, E., Földes, A. and Révész, P.: About the distance between random walkers on some graphs. Period. Math. Hungar. 75 (2017) 36–57.